OpenStudy (anonymous):

which expression makes the equation true? 64h^8=(_)^2 a)32h^8 b)32h^4 c)8h^4 d)8h^8

OpenStudy (cwrw238):

Hint deal with the number 64 and h^8 separately:- 64 = x^2 - what is x ?

OpenStudy (anonymous):

32

OpenStudy (cwrw238):

No - x is the square root of 64 the '^' means to the power 2 or 'squared'

OpenStudy (cwrw238):

* 8 is the square root of 64 because 8*8 = 64 so part of the answer is 8

OpenStudy (anonymous):

8

OpenStudy (cwrw238):

right Now you need to find the square root of h^8

OpenStudy (anonymous):

8h^4

OpenStudy (cwrw238):

tha'ts correct

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

(8h^4)^2 = 8*8* h^4* h^4 = 64 h^8

OpenStudy (cwrw238):

option c

OpenStudy (anonymous):

y^20=(_)^2 would be y^10

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

which pair of numbers id not a factor pair of 40? would it be -8 and 5

OpenStudy (anonymous):

is *

OpenStudy (cwrw238):

yes -8 * 5 = -40

OpenStudy (anonymous):

which pair of numbers is 8 a common factor of? 40 and 48

OpenStudy (anonymous):

would it be 40 and 48

OpenStudy (cwrw238):

yes gotta go